Rocking technologies RT has just developed a sokar panel capable of generating 200% more electricity than amy solar panel currently on the market. As a result, RT is exoected to experience a 15% annual growth rate for the next 5 years. By the end of 5 years, other firms will have developed comparable technology and RT’s growth rate will slow to 5% per year indefinitely. Stockholders require a return of 6% on the stock. The most recent dividend was paid yesterday at $1.55 per share. What are the expected dividends for years 1-5? What is the intrinsic value of the stock?
Answer #1
Expected dividend for next 5 years
Years 1 2 3 4 5 Total
Dividend(expected) 1.7825 2.0499 2.3574 2.7110 3.1176
PVF @ 6% 0.9434 0.8900 0.8396 0.7921 0.7473
PV of dividend 1.6816 1.8244 1.9793 2.1473 2.3297 9.9623
Value of share at the end of year 5 = div at yr 5 end/k-g
= 9.9623/(0.06-0.05)
= 9.9623/0.01
= 996.23
PV of value today = 996.23*PVF (6%,5 years)
= 744.441
Value of share = PV of share price+PV of dividend
= 744.441+9.9623
= $754.40